a body covers first 1/3 part of its journey with velocity= 2 m/s , next1/3 part with velocity of 3m/s and rest of the journey with velocity 6 m/s . Average velocity of body is

Grade:12th pass

3 Answers

Sam
6 years ago
average velocity = displace / time
so here we can assume total displacement covered is S
for 1st  part    time = displacement / velocity   =  (S/3) / 2 
as displacement is S/3 and veocity is 2
 
similarly for 2nd part time = (S/3) / 3
                   3rd part time = (S/3) / 6
 
so total time = (S/6) + (S/9) + (S/18) =  S (3+2+1)     =  S (6)    =   S
                                                                      18                 18          3
 
therefore average velocity =   S      =  S3   =   3m/s
                                             (S/3)       S
Jayashri Walzade
5 years ago
Let total displacement be 3x and it's 1/3rd part is x, x and x
Average velocity =  total displacement /total time
 
For T1(time ) =displacement/time =x/2sec
T2 = x/3 sec
T3 = x/6 sec
Total time = x/2+x/3+x/6
                  =6x/x 
                  =X(total time)
Now, average velocity = 3x/x (as we assume total             displacement =3x above)
                    Ans                =3 m/s
  
                  
       
ankit singh
3 years ago

Explanation:

average velocity = displace / time

so here we can assume total displacement covered is S

for 1st  part    time = displacement / velocity   =  (S/3) / 2  

as displacement is S/3 and veocity is 2

 

similarly for 2nd part time = (S/3) / 3

                  3rd part time = (S/3) / 6

 

so total time = (S/6) + (S/9) + (S/18) =  S (3+2+1)     =  S (6)    =   S

                                                                     18                 18          3

 

therefore average velocity =   S      =  S3   =   3m/s

                                            (S/3)       S